OpenStudy (anonymous):

if, for all x, f'(x)=(x-2)^4(x-1)^3 , it follows that the function f has

myininaya (myininaya):

so you need help integrating then \[\int\limits (x-2)^2(x-1)^3 dx \] Let u=x-1 then du=dx if u=x-1 subtracting one on both sides gives: u-1=x-2 so you have \[\int\limits_{}^{}(u-1)^2u^3 du\] this is just to make the multiplication less harsh we still have to multiply

myininaya (myininaya):

so multiply out (u-1)^2 then multiply each one of those terms by u^3 then integrate term by term

myininaya (myininaya):

oh you aren't asking to find f

myininaya (myininaya):

what are you asking for

OpenStudy (anonymous):

i just started learning calculus and came across this problem on a practice test could u tell me specifically what this is called so i can search it up and learn it

myininaya (myininaya):

what is the question exactly though ?

OpenStudy (anonymous):

i was asking for the relative minimum

OpenStudy (anonymous):

i just dont kno how to approach this

myininaya (myininaya):

well you can find the critical numbers of f by setting f'=0 and also finding where f' dne (we don't have to worry about this case here because f' is a polynomial) f'=0 means we need to solve (x-2)^2(x-1)^3=0 again these x's that satisfy (x-2)^2(x-1)^3=0 are critical numbers

myininaya (myininaya):

the critical numbers is where you can possibly have relative mins or maxes

OpenStudy (anonymous):

oh so i just set x-1=0

myininaya (myininaya):

and x-2=0

myininaya (myininaya):

so you have critical numbers x=1 or x=2 now it is possible to have a rel max or min or neither at the critical numbers

OpenStudy (anonymous):

oh kk and 1 and 2 are the critical number but the relative minimum is 1 since that is the lowest

myininaya (myininaya):

I would rather not find f'' we can use f' and draw a number line in test intervals around these numbers to see if f is decreasing or increasing

myininaya (myininaya):



myininaya (myininaya):

well we don't know that

myininaya (myininaya):

i mean yes 1 is less than 2 but that doesn't mean f(1) is a relative min

myininaya (myininaya):